OpenStudy (anonymous):

2/7+1/2+3/14= answer please

OpenStudy (anonymous):



OpenStudy (lalaly):

\[\frac{2}{7} \times \frac{2}{2} + \frac{1}{2} \times \frac{7}{7} + \frac{3}{14}\]\[= \frac{4}{14} + \frac{7}{14} + \frac{3}{14} = \frac{14}{14} = 1\]

OpenStudy (anonymous):

i know really bad drawing

OpenStudy (anonymous):

(2/7)+(1/2)= find a common denominator 2(2)/7(2)+ 1(7)/2(7)= 11/14+3/14=14/14=1